OpenStudy (anonymous):

A community recreation center is having a funraising event.they are selling products in which they will bring in $500.the entrance fee is $5.00 per person.what equation gives the total cost per person y as a function of the number of people x ,if everyone buys the same amount?

OpenStudy (callisto):

$500. <- per person or total?

OpenStudy (anonymous):

thats the answer?

OpenStudy (callisto):

Nope.. just asking you...

OpenStudy (anonymous):

o idk i typed it in exactly how it says so..

OpenStudy (callisto):

Hmm.. That's what I guess. 500 is the total => each person = 500/x Equation required is y = 500/x + 5